36. The following account balances and information are available for Stevens Inc. at the end of 2019: Overapplied ManufacturiHelp me out
Answer #1
Total Overhead applied 120000 =10000+30000+80000
Overapplied overhead allocated to:
Work in process 5000 =60000*10000/120000
Finished goods 15000 =60000*30000/120000
Cost of goods sold 40000 =60000*80000/120000
Unadjusted Cost of goods sold 200000
Less: Overhead overapplied 40000
Cost of goods sold 160000
Option C $160,000 is correct